# CrystalFlow: A Flow-Based Generative Model for Crystalline Materials

### v3

@@ II Method
This choice of invariant lattice representation $\mathbf{k}$ provides significant flexibility in defining both the prior distribution and the conditional probability path. For simplicity, we adopt a Gaussian prior $q(\mathbf{k}_{0})=\mathcal{N}(\mathbf{k}_{0};\bm{\mu}^{k}_{0},\bm{\sigma}^{k}_{0})$, where the mean and standard deviation are set as $\bm{\mu}^{k}_{0}=\{0,0,0,0,0,1\}$ and $\bm{\sigma}^{k}_{0}=0.1$, respectively. The probability path is defined with a time-dependent mean $\bm{\mu}^{k}_{t}$ determined by linear interpolation between the initial and terminal points, given by $\bm{\mu}^{k}_{t}=t\mathbf{k}_{1}+(1-t)\mathbf{k}_{0}$, and a constant standard deviation $\bm{\sigma}^{k}_{t}=\bm{\sigma}^{k}$. The conditional probability path and vector field are formulated as:
@@
By construction, the conditional velocity field $u^{F}_{t}(\mathbf{F}|z)$ is pairwise invariant under periodic translations, resulting in the conditional probability path $p_{t}^{F}(\mathbf{F}|z)$ being pairwise periodic translation invariant, as demonstrated in Supplementary S1. We set $\bm{\sigma}^{F}=0$ to be consistent with the Rectified Flow framework [52].
@@
where $\lambda_{k}$ and $\lambda_{F}$ are weights for the respective loss terms, $x$ represents the interpolating structure $(\mathbf{k}_{t},\mathbf{F}_{t})$, and $y$ represents the chemical composition $\mathbf{A}$ and external pressure $P$. The vector fields $v_{\theta}^{k}$ and $v_{\theta}^{F}$ are parameterized by neural networks with learnable parameters $\theta$.
@@
where $v_{\theta}^{A}$ is the predicted vector field. The total loss is given by:
@@ Acknowledgements.
The work is supported by the National Key Research and Development Program of China (Grant No. 2022YFA1402304), the National Natural Science Foundation of China (Grants No. 12034009, 12374005, 52288102, 52090024, and T2225013), the Fundamental Research Funds for the Central Universities, the Program for JLU Science and Technology Innovative Research Team. We want to thank “Changchun Computing Center” and “Eco-Innovation Center” for providing inclusive computing power and technical support of MindSpore during the completion of this paper. Part of the calculation was performed in the high-performance computing center of Jilin University. We also want to thank all CALYPSO community contributors for providing the structural data to construct the dataset.
@@
J.L., L.W., Y.W., and Y.M. designed the research. X.L. wrote the code and trained the model. Z.W. and X.L. collected the dataset. X.L. and Z.W. conducted the calculations. X.L., Z.W., J.L., L.W., Y.W., and Y.M. analyzed and interpreted the data, and contributed to the writing of the paper.

### v4

@@ II Method
This choice of invariant lattice representation $\mathbf{k}$ provides significant flexibility in defining both the prior distribution and the conditional probability path. For simplicity, we adopt a Gaussian prior $q(\mathbf{k}_{0})=\mathcal{N}(\mathbf{k}_{0};\bm{\mu}^{k}_{0},\bm{\sigma}^{k}_{0})$, where the mean and standard deviation are set as $\bm{\mu}^{k}_{0}=(0,0,0,0,0,1)$ and $\bm{\sigma}^{k}_{0}=0.1$, respectively. The probability path is defined with a time-dependent mean $\bm{\mu}^{k}_{t}$ determined by linear interpolation between the initial and terminal points, given by $\bm{\mu}^{k}_{t}=t\mathbf{k}_{1}+(1-t)\mathbf{k}_{0}$, and a constant standard deviation $\bm{\sigma}^{k}_{t}=\bm{\sigma}^{k}$. The conditional probability path and vector field are formulated as:
@@
By construction, the conditional vector field $u^{F}_{t}(\mathbf{F}|z)$ is pairwise invariant under periodic translations, resulting in the conditional probability path $p_{t}^{F}(\mathbf{F}|z)$ being pairwise periodic translation invariant, as demonstrated in Supplementary S1. We set $\bm{\sigma}^{F}=0$ to be consistent with the Rectified Flow framework [52].
@@
where $\lambda_{k}$ and $\lambda_{F}$ are weights for the respective loss terms, $x$ represents the interpolating structure $(\mathbf{k}_{t},\mathbf{F}_{t})$, and $y$ represents the chemical composition $\mathbf{A}$ and external pressure $P$. The vector fields $v_{t;\theta}^{k}$ and $v_{t;\theta}^{F}$ are parameterized by neural networks with learnable parameters $\theta$.
@@
where $v_{t;\theta}^{A}$ is the predicted vector field. The total loss is given by:
@@ Acknowledgements.
The work is supported by the National Key Research and Development Program of China (Grant No. 2022YFA1402304), the National Natural Science Foundation of China (Grants No. 12034009, 12374005, 52288102, 52090024, and T2225013), the Fundamental Research Funds for the Central Universities, the Program for JLU Science and Technology Innovative Research Team. We want to thank “Changchun Computing Center” and “Eco-Innovation Center” for providing inclusive computing power and technical support of MindSpore during the completion of this paper and the support provided by MindSpore Community. Part of the calculation was performed in the high-performance computing center of Jilin University. We also want to thank all CALYPSO community contributors for providing the structural data to construct the dataset.
@@
J.L., L.W., Y.W., and Y.M. designed the research. X.L. and Q.W. wrote the code and trained the model. Z.W. and X.L. collected the dataset. X.L., Z.W. and Q.W. conducted the calculations. X.L., Z.W., Q.W., J.L., L.W., Y.W., and Y.M. analyzed and interpreted the data, and contributed to the writing of the paper.